# Cell-cell communication inference and analysis: biological mechanisms, computational approaches, and future opportunities

### v1

@@ 4.4 Optimal Transport Methods
Methods based on optimal transport conceptualize CCC as a resource allocation problem (Figure 3b). These methods typically infer the most probable communication paths between cells by minimizing the transportation cost of signaling molecules across spatial locations. SpaOTsc[RN376SpaOTsc] is a pioneering approach that formulates two sequential optimal transport problems to decode spatial cellular organization and communication. The method first aligns scRNA-seq data (source distribution) with spatial transcriptomics data (target distribution) by computing an optimal transport plan that minimizes the cost based on gene expression dissimilarity. The resulting optimal transport distance quantifies spatial proximity between cell pairs. Subsequently, SpaOTsc infers CCC networks by solving the second optimal transport problem where sender and receiver cells are modeled as source and target distributions, with intercellular distances defining the transport cost. The optimal transport plan directly yields the likelihood of communication for each cell pair, integrating both spatial constraints and expression levels of ligands, receptors, and downstream genes. In addition, SpaOTsc trains a random forest model to predict downstream gene expression, thereby estimating the spatial range over which specific signaling pathways operate. To account for the competition between different ligand and receptor species as well as spatial distances between cells, COMMOT[cang2023screeningCOMMOT] develops a collective optimal transport method to handle complex molecular interactions and spatial constraints. Mathematically, the collective optimal transport problem is formulated as follows:
@@
where $X_{i,k}^{L}$ and $X_{j,l}^{R}$ represent the expression level of ligand $i$ on spot $k$ and the expression level of receptor $j$ on spot $l$. $F$ penalizes the untransported mass $\mu_{i}$ and $\nu_{j}$. $P_{i,j,k,l}$ is the coupling matrix scoring the signaling strength from spot $k$ to spot $l$ through the pair consisting of the ligand $i$ and receptor $j$, while $I$ denote the index set of bindable ligand-receptor pairs. $C_{(i,j)}$ is the cost matrix based on the thresholded distance matrix, designed to impose spatial distance constraints. The collective optimal transport problem determines a collection of optimal transport plans for all pairs of ligand and receptor species that can be coupled simultaneously, thereby enabling the consideration of competition among different species.
Apart from modeling CCC via optimal transport, ScACCorDION[RN427] focuses on characterizing changes in intercellular communication across multiple biological conditions. It represents the communication network between cell types within each sample as a directed weighted graph, and applies optimal transport to compute the Wasserstein distance between these graphs. Additionally, ScACCorDION identifies communication events that differ between sample groups by estimating the barycenters of a collection of CCC networks.
@@ 4.5 Factorization Methods
These methods infer CCC and their latent patterns using matrix or tensor factorization, capable of simultaneously analyzing multiple factors influencing CCC (Figure 3b). Among matrix factorization-based tools[RN419, zhang2023defining, RN424Nico], CSCP[zhang2023defining] aims to reveal communication patterns between cell types at a fine-grained level. It identifies pairs of cell subgroups with strong and similar intercellular crosstalk signals by solving a coupled non-negative matrix factorization problem, which minimizes differences in ligand or receptor expression between cells while maximizing the activity scores of crosstalk signals mediated by ligand–receptor pairs. By integrating scRNA-seq data with spatial transcriptomics data, NiCo[RN424Nico] captures intra-cell type variability through a non-negative matrix factorization framework based on the concept of niche composition, leading to the identification of latent factors for each cell type. These latent factors are then used to infer relevant signaling mediators and cell type pairs with covarying factors, thereby enabling the construction of intercellular interaction networks between cell types.
Methods like DIALOGUE[jerby2022dialogueDIALOGUE], MOFAcellular[flores2023multicellular] and scITD[RN323] uncovers latent multicellular programs from single-cell gene expression data by applying penalized matrix decomposition[witten2009penalized], multi-omics factor analysis (MOFA)[argelaguet2018multi, argelaguet2020mofa] and Tucker tensor decomposition[tucker1966some], respectively. Potential signaling is then identified based on the ligands and receptors enriched in the multicellular programs. Notably, MOFAcellular is a flexible multi-view integration framework, which enables the inclusion of additional tissue-level descriptions in the model, such as celltype compositions, spatial relationships, and CCC scores.
By organizing the inferred CCC between cell types into a tensor, several methods, including scTensor[RN399], Tensor-cell2cell[RN339] and STACCato[Dai20231215571918], have been proposed to uncover CCC patterns as well as their unique combinations of cell types and ligand-receptor pairs by using non-negative Tucker decomposition[kim2007nonnegative], CANDECOMP/PARAFAC (CP) decomposition[carroll1970analysis, harshman1970foundations] and Tucker decomposition, respectively. Tensor-cell2cell is the pioneering method for uncovering context-driven CCC patterns by constructing a 4D-communication tensor. Mathematically, Tensor-cell2cell constructs a $C\times P\times S\times T$ fourth-order tensor $\chi$, where $C$, $P$, $S$, $T$ correspond to the number of contexts, ligand-receptor pairs, sender cells and receiver cells respectively. Tensor factorization is then performed by iteratively optimizing the following objective function until convergence:
@@
where $\otimes$ denotes the outer product and $c^{r}$, $p^{r}$, $s^{r}$ and $t^{r}$ are vectors of the factor $r$, each containing the loadings of the respective elements along a specific dimension of the tensor, enabling the further identification and interpretation of context-dependent communication patterns. Following this idea, STACCato identifies condition-related CCC events while also explicitly adjusting for sample-level confounding variables (e.g., batch, age and gender). More recently, Tensor-cell2cell v2[Armingol20221102514917] performs integrative analysis of protein- and metabo-litemediated CCC by using coupled tensor component analysis, thereby facilitating the joint interpretation of CCC patterns across multiple modalities.
@@ 5 Intuitive Visualization and Systems Analysis of Cell-Cell Communication
For network visualization, tools employ diverse graphical representations: CellChat[RN342, jin2025cellchat] utilizes circle plots, hierarchical diagrams, chord diagrams, heatmaps, bubble plots, and word clouds to effectively illustrate intercellular communication patterns and highlight signaling variations across cell types (Figure 6a). Similarly, CellCall[RN268] employs alluvial plots to trace cascading relationships from ligands through receptors to downstream target genes, while NicheNet[RN349, sang2025unraveling] provides heatmaps displaying ligand-target regulatory potential and predicted ligand activity (Figure 6b). In spatial contexts, CellChat visualizes communication links directly on tissue architecture, and COMMOT[cang2023screeningCOMMOT] maps signaling activity to individual cellular locations (Figure 6c).
@@
Spatial analysis reveals distinctive organizational patterns of CCC (Figure 6e). COMMOT interpolates communication patterns into vector fields to visualize spatial directionality of signal transmission and reception. CytoSignal[RN410CytoSignal] introduces signaling velocity by leveraging RNA velocity of ligands and receptors, enabling identification and visualization of dynamic signal intensity changes across tissue space. Meanwhile, MintFlow[Akbarnejad2025] and GITIII[Xiao20240821608964GITIII] generate microenvironment-induced embeddings that enable fine-grained cell type separation based on communication-driven cellular states.
@@
Comparative analysis across biological conditions detects complex CCC changes under varying contexts. CellChat identifies altered signaling pathways and ligand-receptor pairs by analyzing differences in network architecture, information flow, and through joint manifold learning of intrinsic CCC structures (Figure 6f). Beyond the pairwise comparison, Tensor-cell2cell[RN339] employs tensor decomposition to uncover context-driven communication patterns with unique combinations of cell types and ligand-receptor pairs, visualizing context loadings, ligand-receptor pair loadings, and sender/receiver cell loadings to facilitate pattern interpretation across diverse conditions (Figure 6f). STACCato[Dai20231215571918] identifies condition-related CCC while adjusting for sample-level variables (e.g., batch, age, gender) using supervised tensor regression. MultiNicheNet[RN336] addresses differential CCC analysis in complex multifactorial designs, accommodating inter-sample heterogeneity while correcting for batch effects and covariates through ranking-based prioritization integrated into an aggregated score that highlights significantly altered interactions for downstream interpretation (Figure 6f).
@@ 6.1 Resources, Validation and Benchmarking
The development of CCC analysis tools relies on diverse prior knowledge resources and methodological frameworks. Systematic comparisons reveal that choices of resources and methods substantially impact inference results, thereby influencing downstream biological interpretations[LIANA].
@@
Accurate recapitulation of molecular interactions is essential for biologically meaningful CCC prediction. However, ligand–receptor databases vary considerably in content—including multisubunit complex annotation, functional classifications (e.g., secreting signaling or contact-dependent signaling, molecular function and subcellular location), inclusion of intracellular signaling events, and the total number of curated interactions. Detailed comparisons of existing databases are available in a previous review[cesaro2025advances]. In addition to the protein-mediated ligand-receptor databases, resources are emerging for metabolite–protein interactions (e.g., MACC[gao2024macc], MetalinksDB[farr2024metalinksdb], MRCLinkdb[zhang2024predicting]) and EV-mediated interactions (e.g., EV-COMM[chen2024ev], miRTalkDB[shao2025extracellular]). Since most tools rely on their own or referenced databases—and results are highly sensitive to database quality—researchers may draw different conclusions using different methods. A unified, well-annotated, high-quality interaction repository is urgently needed to standardize and enhance the reliability of CCC inference.
@@
Figure 7: Challenges and opportunities in CCC analysis. (a) Validation and benchmarking challenges. Inferred CCC results are influenced by the choice of prior knowledge resources. However, the absence of experimental ground truth and standardized prior knowledge resources hinders systematic tool evaluation. (b) Modeling CCC with precise mechanisms and other modalities. Most CCC studies focus on protein-mediated ligand–receptor interactions, overlooking non-protein signaling and complex regulatory mechanisms. Advances in multi-omics technologies and de novo network reconstruction now enable more comprehensive CCC inference and the discovery of previously uncharacterized CCC. (c) Spatiotemporal and cross-condition analysis. Integrating spatial and temporal context, along with comparative studies across conditions, can reveal deeper biological insights—yet these approaches face considerable technical and analytical hurdles. (d) Clinical and phenotypic integration. Linking phenotypic and clinical data to CCC helps identify key CCC in disease progression, supporting the development of precision medicine. (e) Mechanistic-informed modeling. Further refinement to incorporate communication more explicitly will not only improve the accuracy of the learned cellular dynamics but also offer new insights into temporal CCC dynamics and cellular responses to altered cellular communication.Mathematical models of signaling mechanisms allow researchers to probe regulatory dynamics under experimentally inaccessible conditions.. (f) Engineering and intervention. Emerging computational approaches—including in silico perturbation, foundation models, and virtual cells—open new avenues for predictive and engineering-based CCC research, though model interpretability remains a critical challenge. Created in https://BioRender.com
@@
A major challenge remains the scarcity of experimental ground truth for evaluating computational predictions (Figure 7a). Recent experimental advances—including barcode-based (BRICseq[BRICseq2020], RABID-seq[RABID-seq2021]) and droplet-based (ProximID[ProximID2018], PIC-seq[PIC-seq2020]) technologies—now enable tracing of sender–receiver contacts at single-cell resolution, thereby facilitating validation of predicted CCC. For instance, PIC-seq identifies context-specific cell interactions and differentially regulated genes, while SPEAC-seq[wheeler2023droplet] captures soluble ligands to link sender-cell perturbations with receiver-cell responses. Although these methods improve measurement throughput and accuracy, most remain limited to direct cellular contacts and cannot simultaneously profile multiple ligand–receptor interactions or cell pairs. Further innovation in experimental techniques and curation of ground-truth datasets are essential for systematic validation of computational predictions.
@@
Computational benchmarking using simulated data with known ground truth has been attempted in several studies[li2025scmultisim, RN454stMLnet, cang2023screeningCOMMOT, tanevski2022explainable]. However, accurately simulating CCC remains challenging due to the complexity and redundancy of signaling in multicellular systems. Systematic tool assessment is further complicated by the diversity of underlying databases, biological assumptions, data compatibility, and parameter settings. Although recent benchmarking efforts have established more systematic evaluation frameworks—particularly for scRNA-seq-based methods[RN305, benchmarkCCCWang]—development of objective and standardized quantitative metrics remains crucial. Accessible and unified platforms such as LIANA+[dimitrov2024liana+] are emerging to address this need. With over 50 computational methods now available for spatial transcriptomics data, systematic benchmarking of these spatially aware tools is urgently required to guide future methodological development.
@@ 6.2 Comprehensive modeling of CCC with precise mechanisms
While most CCC inference methods focus on protein-mediated LRIs, non-protein signaling mechanisms, including those mediated by small molecules such as metabolites and neurotransmitters, also play vital roles in biological systems and require systematic investigation (Figure 7b). For instance, profiling enzymes and transporters involved in non-protein signaling enables inference of metabolite-mediated CCC[zheng2022mebocost] or neuron–neuron communication[zhao2023inferring, jakobsson2021scconnect] from scRNA-seq data. A central challenge in metabolite-mediated CCC analysis lies in accurately estimating metabolic activity. Recently, a scalable tool called scCellFie[Armingol20250509653038] has been developed to infer metabolic activity per single cell or spatial spot from transcriptomic data. Non-protein signaling molecules exhibit greater diversity and more complex regulatory mechanisms than proteins, necessitating careful consideration in modeling.
@@
Beyond metabolite signaling, EVs are increasingly recognized as key mediators of CCC in both prokaryotic and eukaryotic organisms, with considerable clinical potential as biomarkers and drug delivery vehicles[van2022challenges]. Recent years have seen the emergence of databases, technologies, and computational tools to elucidate EV-mediated communication dynamics. For example, EV-COMM[chen2024ev] compiles literature-curated EV-mediated interactions across cells and species, linking EV cargo to downstream pathways and functional outcomes. Similarly, miRTalkDB[shao2025extracellular] catalogs experimentally supported EV-derived miRNAs and their target interactions in humans, mice, and rats, providing a foundation for inferring EV–miRNA-mediated CCC. Functional RNAs within EVs—such as mRNAs and miRNAs—can modulate recipient cell behavior and are critical to intercellular signaling. However, standard scRNA-seq protocols are optimized for cellular RNAs and often fail to capture EV RNAs effectively[miceli2024extracellular]. Bridging this technical gap requires specialized experimental and computational workflows that account for the low abundance, heterogeneity, and fragmented nature of EV transcripts. Recent advances enable characterization of EV transcriptomic features, estimation of cellular secretion activity, and related analyses[luo2022transcriptomic, he2024sevtras, Zhao20250624660984]. For instance, SEVtras[he2024sevtras] identifies droplets containing small extracellular vesicles and infers a vesicle signal score to estimate single-cell secretion activity from scRNA-seq data. More recently, miRTalk[shao2025extracellular] has emerged as the first method to systematically infer EV-derived miRNA-mediated CCC by integrating module scores for EV biogenesis, secretion signatures, RNA-induced silencing complex–related genes, and miRNA–target expression. Despite these advances, the mechanisms and applications of EV-mediated CCC remain incompletely understood. Further investigation—particularly into vesicle heterogeneity in spatial contexts—will deepen insights into EV-driven communication.
@@
Most current CCC studies rely on mRNA expression as a proxy for protein abundance, an assumption often undermined by post-transcriptional regulation and complex protein trafficking processes. The advent of single-cell and spatial multi-omics technologies—including proteomics[stoeckius2017simultaneous, peterson2017multiplexed], epigenomics[cusanovich2015multiplex], and metabolomics[lan2024single, mao2025single]—enables more comprehensive and granular characterization of CCC. Integrating single-cell ATAC-seq data, for example, can illuminate transcriptional regulation dynamics[jin2020scai]. Moreover, intracellular post-translational modifications (PTMs) play central roles in signal integration and gene regulation. Emerging multimodal profiling technologies such as Phospho-seq[blair2025phospho] and SIGNAL-seq[Opzoomer20240223581433] allow measurement of PTMs, offering deep insights into cell signaling and transcriptional responses. SIGNAL-seq, in particular, enables simultaneous profiling of mRNA ligand–receptor pairs, intracellular PTMs, and transcriptional states in thousands of single cells. Expanding analytical tools to incorporate such multi-omics data will improve both inference and validation of CCC and support more effective benchmarking.
@@
Advances in computational methods now enable more accurate modeling of complex CCC mechanisms, including signaling competition and cooperation, feedback regulation, and higher-order interactions. For example, COMMOT[cang2023screeningCOMMOT] accounts for competition among ligands and receptors; SigXTalk[hou2025dissecting] quantifies signaling crosstalk fidelity and specificity using scRNA-seq data; and LRLoop[RN288] identifies intercellular feedback loops. A remaining challenge is the integration of diverse mechanisms to evaluate intercellular communication potential at a global scale. With continued research, dynamic and mechanistic models of CCC are expected to become feasible, promoting a systems-level understanding of cellular behavior.
@@
Beyond conventional CCC inference, new computational approaches identify external signals not captured in the data, detect hierarchical communication structures, and reveal intercellular signaling flows. exFINDER[RN281] infers external signals and their downstream networks by leveraging prior knowledge of ligand–receptor–transcription factor pathways and scRNA-seq data, addressing signals originating outside the profiled cellular population. CrossChat[RN409] detects hierarchical structures in CCC through two complementary strategies: CrossChatH performs multi-resolution cell clustering to reveal global communicative hierarchies, and CrossChatT identifies local tree structures among ligands and receptors based on inclusive or disjoint expression patterns. To identify LRIs where intracellular processes mediate signal inflow and trigger outflow of other intercellular signals, FlowSig[RN415FlowSig] uses graphical causal modeling with conditional independence tests to infer a completed partially directed acyclic graph (CPDAG) representing intercellular flows.

### v3

@@ 4.4 Optimal Transport Methods
Methods based on optimal transport conceptualize CCC as a resource allocation problem (Figure 3b). These methods typically infer the most probable communication paths between cells by minimizing the transportation cost of signaling molecules across spatial locations. SpaOTsc[28] is a pioneering approach that formulates two sequential optimal transport problems to decode spatial cellular organization and communication. The method first aligns scRNA-seq data (source distribution) with spatial transcriptomics data (target distribution) by computing an optimal transport plan that minimizes the cost based on gene expression dissimilarity. The resulting optimal transport distance quantifies spatial proximity between cell pairs. Subsequently, SpaOTsc infers CCC networks by solving the second optimal transport problem where sender and receiver cells are modeled as source and target distributions, with intercellular distances defining the transport cost. The optimal transport plan directly yields the likelihood of communication for each cell pair, integrating both spatial constraints and expression levels of ligands, receptors, and downstream genes. In addition, SpaOTsc trains a random forest model to predict downstream gene expression, thereby estimating the spatial range over which specific signaling pathways operate. To account for the competition between different ligand and receptor species as well as spatial distances between cells, COMMOT[29] develops a collective optimal transport method to handle complex molecular interactions and spatial constraints. Mathematically, the collective optimal transport problem is formulated as follows:
@@
where $X_{i,k}^{L}$ and $X_{j,l}^{R}$ represent the expression level of ligand $i$ on spot $k$ and the expression level of receptor $j$ on spot $l$. $F$ penalizes the untransported mass $\mu_{i}$ and $\nu_{j}$. $P_{i,j,k,l}$ is the coupling matrix scoring the signaling strength from spot $k$ to spot $l$ through the pair consisting of the ligand $i$ and receptor $j$, while $I$ denotes the index set of bindable ligand-receptor pairs. $C_{(i,j)}$ is the cost matrix based on the thresholded distance matrix, designed to impose spatial distance constraints. The collective optimal transport problem determines a collection of optimal transport plans for all pairs of ligand and receptor species that can be coupled simultaneously, thereby enabling the consideration of competition among different species.
Apart from modeling CCC via optimal transport, ScACCorDION[130] focuses on characterizing changes in intercellular communication across multiple biological conditions. It represents the communication network between cell types within each sample as a directed weighted graph, and applies optimal transport to compute the Wasserstein distance between these graphs. Additionally, ScACCorDION identifies communication events that differ between sample groups by estimating the barycenters of a collection of CCC networks.
@@ 4.5 Factorization Methods
These methods infer CCC and their latent patterns using matrix or tensor factorization, capable of simultaneously analyzing multiple factors influencing CCC (Figure 3b). Among matrix factorization-based tools[115, 198, 1], CSCP[198] aims to reveal communication patterns between cell types at a fine-grained level. It identifies pairs of cell subgroups with strong and similar intercellular crosstalk signals by solving a coupled non-negative matrix factorization problem, which minimizes differences in ligand or receptor expression between cells while maximizing the activity scores of crosstalk signals mediated by ligand–receptor pairs. By integrating scRNA-seq data with spatial transcriptomics data, NiCo[1] captures intra-cell type variability through a non-negative matrix factorization framework based on the concept of niche composition, leading to the identification of latent factors for each cell type. These latent factors are then used to infer relevant signaling mediators and cell type pairs with covarying factors, thereby enabling the construction of intercellular interaction networks between cell types.
Methods like DIALOGUE[82], MOFAcellular[57] and scITD[125] uncover latent multicellular programs from single-cell gene expression data by applying penalized matrix decomposition[183], multi-omics factor analysis (MOFA)[7, 6] and Tucker tensor decomposition[167], respectively. Potential signaling is then identified based on the ligands and receptors enriched in the multicellular programs. Notably, MOFAcellular is a flexible multi-view integration framework, which enables the inclusion of additional tissue-level descriptions in the model, such as cell type compositions, spatial relationships, and CCC scores.
By organizing the inferred CCC between cell types into a tensor, several methods, including scTensor[166], Tensor-cell2cell[10] and STACCato[43], have been proposed to uncover CCC patterns as well as their unique combinations of cell types and ligand-receptor pairs by using non-negative Tucker decomposition[95], CANDECOMP/PARAFAC (CP) decomposition[30, 68] and Tucker decomposition, respectively. Tensor-cell2cell is the pioneering method for uncovering context-driven CCC patterns by constructing a 4D-communication tensor. Mathematically, Tensor-cell2cell constructs a $C\times P\times S\times T$ fourth-order tensor $\chi$, where $C$, $P$, $S$, $T$ correspond to the number of contexts, ligand-receptor pairs, sender cells and receiver cells respectively. Tensor factorization is then performed by iteratively optimizing the following objective function until convergence:
@@
where $\otimes$ denotes the outer product and $c^{r}$, $p^{r}$, $s^{r}$ and $t^{r}$ are vectors of the factor $r$, each containing the loadings of the respective elements along a specific dimension of the tensor, enabling the further identification and interpretation of context-dependent communication patterns. Following this idea, STACCato identifies condition-related CCC events while also explicitly adjusting for sample-level confounding variables (e.g., batch, age and gender). More recently, Tensor-cell2cell v2[12] performs integrative analysis of protein- and metabo-litemediated CCC by using coupled tensor component analysis, thereby facilitating the joint interpretation of CCC patterns across multiple modalities.
@@ 5 Intuitive Visualization and Systems Analysis of Cell-Cell Communication
For network visualization, tools employ diverse graphical representations: CellChat[88, 89] utilizes circle plots, hierarchical diagrams, chord diagrams, heatmaps, bubble plots, and word clouds to effectively illustrate intercellular communication patterns and highlight signaling variations across cell types (Figure 6a). Similarly, CellCall[201] employs alluvial plots to trace cascading relationships from ligands through receptors to downstream target genes, while NicheNet[24, 149] provides heatmaps displaying ligand-target regulatory potential and predicted ligand activity (Figure 6b). In spatial contexts, CellChat visualizes communication links directly on tissue architecture, and COMMOT[29] maps signaling activity to individual cellular locations (Figure 6c).
@@
Spatial analysis reveals distinctive organizational patterns of CCC (Figure 6e). COMMOT interpolates communication patterns into vector fields to visualize spatial directionality of signal transmission and reception. CytoSignal[110] introduces signaling velocity by leveraging RNA velocity of ligands and receptors, enabling identification and visualization of dynamic signal intensity changes across tissue space. Meanwhile, MintFlow[2] and GITIII[187] generate microenvironment-induced embeddings that enable fine-grained cell type separation based on communication-driven cellular states.
@@
Comparative analysis across biological conditions detects complex CCC changes under varying contexts. CellChat identifies altered signaling pathways and ligand-receptor pairs by analyzing differences in network architecture, information flow, and through joint manifold learning of intrinsic CCC structures (Figure 6f). Beyond the pairwise comparison, Tensor-cell2cell[10] employs tensor decomposition to uncover context-driven communication patterns with unique combinations of cell types and ligand-receptor pairs, visualizing context loadings, ligand-receptor pair loadings, and sender/receiver cell loadings to facilitate pattern interpretation across diverse conditions (Figure 6f). STACCato[43] identifies condition-related CCC while adjusting for sample-level variables (e.g., batch, age, gender) using supervised tensor regression. MultiNicheNet[23] addresses differential CCC analysis in complex multifactorial designs, accommodating inter-sample heterogeneity while correcting for batch effects and covariates through ranking-based prioritization integrated into an aggregated score that highlights significantly altered interactions for downstream interpretation (Figure 6f).
@@ 6.1 Resources, Validation and Benchmarking
The development of CCC analysis tools relies on diverse prior knowledge resources and methodological frameworks. Systematic comparisons reveal that choices of resources and methods substantially impact inference results, thereby influencing downstream biological interpretations[47].
@@
Accurate recapitulation of molecular interactions is essential for biologically meaningful CCC prediction. However, ligand–receptor databases vary considerably in content—including multisubunit complex annotation, functional classifications (e.g., secreted signaling or contact-dependent signaling, molecular function and subcellular location), inclusion of intracellular signaling events, and the total number of curated interactions. Detailed comparisons of existing databases are available in a previous review[31]. In addition to the protein-mediated ligand-receptor databases, resources are emerging for metabolite–protein interactions (e.g., MACC[59], MetalinksDB[53], MRCLinkdb[202]) and EV-mediated interactions (e.g., EV-COMM[32], miRTalkDB[154]). Since most tools rely on their own or referenced databases—and results are highly sensitive to database quality—researchers may draw different conclusions using different methods. A unified, well-annotated, high-quality interaction repository is urgently needed to standardize and enhance the reliability of CCC inference.
@@
Figure 7: Challenges and opportunities in CCC analysis. (a) Validation and benchmarking challenges. Inferred CCC results are influenced by the choice of prior knowledge resources. However, the absence of experimental ground truth and standardized prior knowledge resources hinders systematic tool evaluation. (b) Modeling CCC with precise mechanisms and other modalities. Most CCC studies focus on protein-mediated ligand–receptor interactions, overlooking non-protein signaling and complex regulatory mechanisms. Advances in multi-omics technologies and de novo network reconstruction now enable more comprehensive CCC inference and the discovery of previously uncharacterized CCC. (c) Spatiotemporal and cross-condition analysis. Integrating spatial and temporal context, along with comparative studies across conditions, can reveal deeper biological insights—yet these approaches face considerable technical and analytical hurdles. (d) Clinical and phenotypic integration. Linking phenotypic and clinical data to CCC helps identify key CCC in disease progression, supporting the development of precision medicine. (e) Mechanistic-informed modeling. Further refinement to incorporate communication more explicitly will not only improve the accuracy of the learned cellular dynamics but also offer new insights into temporal CCC dynamics and cellular responses to altered cellular communication.Mathematical models of signaling mechanisms allow researchers to probe regulatory dynamics under experimentally inaccessible conditions. (f) Engineering and intervention. Emerging computational approaches—including in silico perturbation, foundation models, and virtual cells—open new avenues for predictive and engineering-based CCC research, though model interpretability remains a critical challenge. Created in BioRender. S. Jin, 2026, https://BioRender.com/y1iwqii.
@@
A major challenge remains the scarcity of experimental ground truth for evaluating computational predictions (Figure 7a). Recent experimental advances—including barcode-based (BRICseq[80], RABID-seq[38]) and droplet-based (ProximID[21], PIC-seq[63]) technologies—now enable tracing of sender–receiver contacts at single-cell resolution, thereby facilitating validation of predicted CCC. For instance, PIC-seq identifies context-specific cell interactions and differentially regulated genes, while SPEAC-seq[181] captures soluble ligands to link sender-cell perturbations with receiver-cell responses. Although these methods improve measurement throughput and accuracy, most remain limited to direct cellular contacts and cannot simultaneously profile multiple ligand–receptor interactions or cell pairs. Further innovation in experimental techniques and curation of ground-truth datasets are essential for systematic validation of computational predictions.
@@
Computational benchmarking using simulated data with known ground truth has been attempted in several studies[104, 191, 29, 162]. However, accurately simulating CCC remains challenging due to the complexity and redundancy of signaling in multicellular systems. Systematic tool assessment is further complicated by the diversity of underlying databases, biological assumptions, data compatibility, and parameter settings. Although recent benchmarking efforts have established more systematic evaluation frameworks—particularly for scRNA-seq-based methods[119, 116]—development of objective and standardized quantitative metrics remains crucial. Accessible and unified platforms such as LIANA+[46] are emerging to address this need. With over 50 computational methods now available for spatial transcriptomics data, systematic benchmarking of these spatially aware tools is urgently required to guide future methodological development.
@@ 6.2 Comprehensive modeling of CCC with precise mechanisms
While most CCC inference methods focus on protein-mediated LRIs, non-protein signaling mechanisms, including those mediated by small molecules such as metabolites and neurotransmitters, also play vital roles in biological systems and require systematic investigation (Figure 7b). For instance, profiling enzymes and transporters involved in non-protein signaling enables inference of metabolite-mediated CCC[208] or neuron–neuron communication[206, 81] from scRNA-seq data. A central challenge in metabolite-mediated CCC analysis lies in accurately estimating metabolic activity. Recently, a scalable tool called scCellFie[8] has been developed to infer metabolic activity per single cell or spatial spot from transcriptomic data. Non-protein signaling molecules exhibit greater diversity and more complex regulatory mechanisms than proteins, necessitating careful consideration in modeling.
@@
Beyond metabolite signaling, EVs are increasingly recognized as key mediators of CCC in both prokaryotic and eukaryotic organisms, with considerable clinical potential as biomarkers and drug delivery vehicles[169]. Recent years have seen the emergence of databases, technologies, and computational tools to elucidate EV-mediated communication dynamics. For example, EV-COMM[32] compiles literature-curated EV-mediated interactions across cells and species, linking EV cargo to downstream pathways and functional outcomes. Similarly, miRTalkDB[154] catalogs experimentally supported EV-derived miRNAs and their target interactions in humans, mice, and rats, providing a foundation for inferring EV–miRNA-mediated CCC. Functional RNAs within EVs—such as mRNAs and miRNAs—can modulate recipient cell behavior and are critical to intercellular signaling. However, standard scRNA-seq protocols are optimized for cellular RNAs and often fail to capture EV RNAs effectively[124]. Bridging this technical gap requires specialized experimental and computational workflows that account for the low abundance, heterogeneity, and fragmented nature of EV transcripts. Recent advances enable characterization of EV transcriptomic features, estimation of cellular secretion activity, and related analyses[120, 71, 207]. For instance, SEVtras[71] identifies droplets containing small extracellular vesicles and infers a vesicle signal score to estimate single-cell secretion activity from scRNA-seq data. More recently, miRTalk[154] has emerged as the first method to systematically infer EV-derived miRNA-mediated CCC by integrating module scores for EV biogenesis, secretion signatures, RNA-induced silencing complex–related genes, and miRNA–target expression. Despite these advances, the mechanisms and applications of EV-mediated CCC remain incompletely understood. Further investigation—particularly into vesicle heterogeneity in spatial contexts—will deepen insights into EV-driven communication.
@@
Most current CCC studies rely on mRNA expression as a proxy for protein abundance, an assumption often undermined by post-transcriptional regulation and complex protein trafficking processes. The advent of single-cell and spatial multi-omics technologies—including proteomics[159, 141], epigenomics[42], and metabolomics[100, 121]—enables more comprehensive and granular characterization of CCC. Integrating single-cell ATAC-seq data, for example, can illuminate transcriptional regulation dynamics[90]. Moreover, intracellular post-translational modifications (PTMs) play central roles in signal integration and gene regulation. Emerging multimodal profiling technologies such as Phospho-seq[20] and SIGNAL-seq[134] allow measurement of PTMs, offering deep insights into cell signaling and transcriptional responses. SIGNAL-seq, in particular, enables simultaneous profiling of mRNA ligand–receptor pairs, intracellular PTMs, and transcriptional states in thousands of single cells. Expanding analytical tools to incorporate such multi-omics data will improve both inference and validation of CCC and support more effective benchmarking.
@@
Advances in computational methods now enable more accurate modeling of complex CCC mechanisms, including signaling competition and cooperation, feedback regulation, and higher-order interactions. For example, COMMOT[29] accounts for competition among ligands and receptors; SigXTalk[76] quantifies signaling crosstalk fidelity and specificity using scRNA-seq data; and LRLoop[189] identifies intercellular feedback loops. A remaining challenge is the integration of diverse mechanisms to evaluate intercellular communication potential at a global scale. With continued research, dynamic and mechanistic models of CCC are expected to become feasible, promoting a systems-level understanding of cellular behavior.
@@
Beyond conventional CCC inference, new computational approaches identify external signals not captured in the data, detect hierarchical communication structures, and reveal intercellular signaling flows. exFINDER[70] infers external signals and their downstream networks by leveraging prior knowledge of ligand–receptor–transcription factor pathways and scRNA-seq data, addressing signals originating outside the profiled cellular population. CrossChat[180] detects hierarchical structures in CCC through two complementary strategies: CrossChatH performs multi-resolution cell clustering to reveal global communicative hierarchies, and CrossChatT identifies local tree structures among ligands and receptors based on inclusive or disjoint expression patterns. To identify LRIs where intracellular processes mediate signal inflow and trigger outflow of other intercellular signals, FlowSig[4] uses graphical causal modeling with conditional independence tests to infer a completed partially directed acyclic graph (CPDAG) representing intercellular flows.